OpenStudy (constanceleak):

Which is an equation of the given line in standard form? A. –8x + 9y = 23 B. –8x + 9y = –23 C. –8x + 7y = 25 D. –9x + 8y = –23

OpenStudy (constanceleak):

pls explain how

OpenStudy (anonymous):

Use algebra to solve for Y

OpenStudy (anonymous):

Y= mx+b

OpenStudy (constanceleak):

umm, but I'm supposed to put the thing in standard form

OpenStudy (constanceleak):

Btw is 2/9 the slope

OpenStudy (anonymous):

My bad i didn't read correctly nor opened your link. the answer is (A) use formula y-y1=m(x-x1)

OpenStudy (constanceleak):

thx!